OpenStudy (anonymous):

For two independent events A and B, P(A or B) = 0.70, P(A) = 0.50, and P(B) = 0.45. Find P(A and B)

OpenStudy (ash2326):

We know that \[P(A\ or\ B)=P(A)+P(B)-P(A\ and\ B)\] so \[P(A\ and\ B)=P(A)+P(B)-P(A\ or\ B)\] I think you can find now?

OpenStudy (anonymous):

so i wuld do 50%+45%-70%??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well maybe if the events are independent, then \[P(A\cap B)=P(A)\times P(B)\]

OpenStudy (anonymous):

so 50% *45%

OpenStudy (anonymous):

if they are independent you should get the same answer in either case, so i should not have said "no" but rather should have said you can use either method

OpenStudy (anonymous):

so 22.5%

OpenStudy (anonymous):

\[.5+.45-.7=.25\] and also \[.5\times .45=.225\] so in fact since they are not the same answer the events are NOT independent and the question is wrong!!

OpenStudy (anonymous):

its not a true or false question

OpenStudy (anonymous):

hmmm where did this question come from? it is always true that \[P(A\cap B)=P(A)+P(B)-P(A\cup B)\] but if the events are independent then it is also true that \[P(A\cap B)=P(A)\times P(B)\] so there is some mistake here

OpenStudy (anonymous):

i copyed and pastened it so it wasnt my typo

OpenStudy (anonymous):

i believe you is it from a book or a worksheet or what? if you want, you should alert your teacher to the fact that the events here cannot be independent with the probabilities given

OpenStudy (anonymous):

because in this case \[.5+.45-.7\neq .5\times .45\] so the events are definitely not independent

OpenStudy (anonymous):

i dont exactl yhave a techer

OpenStudy (anonymous):

where does the question come from?

OpenStudy (anonymous):

my online schol

OpenStudy (anonymous):

write them and tell them there is a mistake in the question if they remove the word "independent" then it would be find, use method one and get \(.45-.5-.7=.25\) but you cannot just claim two events are independent. you have to show that they are

OpenStudy (anonymous):

*fine

OpenStudy (anonymous):

i suspect that the questions are written without much care, or expertise on the part of the writers. someone just wrote it down without thinking